GOLF SWING MOTION ANALYSIS: AN EXPERIMENT ON THE USE OF VERBAL ANALYSIS IN STATISTICAL REASONING

H. AKAIKE

Emeritus Professor, The Institute of Statistical Mathematics

(Received August 20, 2000; revised August 30, 2000)

Abstract. The complexity of human body demands statistical reasoning to perform an analysis of the golf swing. Such reasoning depends on the use of the Informational Data Set, composed of the related objective knowledge, empirical findings, and observational data. Since the image can be described only verbally and the Informational Data Set requires verbal handling the analysis is performed by constructing a verbal model of the swing motion. The analysis reveals two prototypes of swing motion that can produce an image of actual swing motion. It also produces structural explanations of some of the findings by Ben Hogan. The cusp appearing in the locus of the grip is pointed out as the key to the successful realization of the golf swing. It is tacitly assumed that the golfer is right-handed. The possibility of drastic expansion of the realm of statistics by the use of verbal modeling is pointed out.

Key words and phrases: Verbal modeling, Informational Data Set, golf swing, statistical reasoning.